# DiarizationLM: Speaker Diarization Post-Processing with Large Language Models

### 16 Jan 2024 v2

@@ 1 Introduction
In many cases, such errors can usually be fixed by leveraging semantic information from the ASR transcripts. Take Fig. 1 as an example, simply by looking at the textual transcript “good morning how are you”, if we know it consists of two speakers, we can easily tell which word comes from which speaker confidently without using any speaker diarization system. In practice, diarization errors can be much more complicated than the simple example in Fig. 1. To handle such cases, we propose DiarizationLM, a framework to post-process the orchestrated ASR and speaker diarization outputs with a large language model (LLM). While the experiments performed in this paper mainly focus on reducing word diarization errors using LLM, we also show examples how LLM can be used for other purposes such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module in Appendix A.
@@ 4.3 Models
For the ASR model in Fig. 2, we use a universal speech model (USM)  [44] with 600 million parameters trained with the RNN-T loss [45]. For the speaker diarization model in Fig. 2, we use the turn-to-diarize system [7] with a multi-stage clustering setup [17] in our experiments. The number of speakers is unknown (from $1$ to $\infty$) to the speaker diarization system in all of our experiments. However, we would like to point out that the proposed framework is very generic and should work with other ASR or speaker diarization systems as well, such as variants of end-to-end speaker diarization models [20, 21, 22, 23].
@@
For the LLM in Fig. 2, we take a PaLM 2-S model (“text-bison” model in Google Cloud API) [46] as our foundation model, and finetune it on the dataset described in Section 4.1 with data processing steps described in Section 3.5. This model uses a sentence piece model (SPM) of 256k tokens as its tokenizer [47]. During finetuning, we limit the LLM input size by 4,096 tokens, and segment our training and testing data accordingly.
@@
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix " --> " (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix " [eod]" (short for “end of document”; note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
4.4 Experimental results
@@
In Table 1, we show the evaluation results of the USM + turn-to-diarize baseline together with the outputs post-processed by DiarizationLM with three different finetuning flavors. On both datasets, we observe big improvement of both WDER and cpWER with any of the three finetuning flavors. Interesting, the biggest improvement is observed with the hyp2ora flavor, while the smallest improvement is observed with the deg2ref flavor. Specifically for hyp2ora, we see a rel. 55.5% improvement of WDER after post-processing with DiarizationLM on the Fisher testing set. Even if we did not use any Callhome data during the LLM finetuning, we see a rel. 44.9% improvement of WDER on the Callhome testing set. The WER of the USM on the two testing sets are relatively high due to domain mismatch and suboptimal annotation quality of the ground truth. However, this also demonstrated that the DiarizationLM solution provides consistent quality gains even with out-of-domain ASR and speaker diarization models.
@@
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM with three different finetuning flavors. All numbers are percentages.
Testing set USM + turn-to-diarize DiarizationLM
hyp2ora flavor deg2ref flavor mixed flavor
WER WDER cpWER WDER cpWER WDER cpWER WDER cpWER
Fisher 15.48 5.32 21.19 2.37 16.93 3.94 18.55 2.41 16.94
Callhome 15.36 7.72 24.39 4.25 20.22 5.33 21.47 4.76 20.84
Table 2: Evaluation results of the turn-to-diarize baseline system with reference ASR transcript (assuming WER=0%) and the results post-processed by DiarizationLM. All numbers are percentages.
Testing set Reference + turn-to-diarize DiarizationLM
@@
Fisher 2.81 5.19 1.18 2.21
Callhome 3.74 6.82 1.49 2.66
@@ 5 Discussion and future work
Another research direction would be to compare different LLMs, in different size variants on the speaker diarization task. It would also be interesting to reproduce the experiments with other speaker diarization systems such as EEND [20] or WEEND [34].

### 06 Feb 2024 v4

@@ 1 Introduction
In many cases, such errors can usually be fixed by leveraging semantic information from the ASR transcripts. Take Fig. 1 as an example, simply by looking at the textual transcript “good morning how are you”, if we know it consists of two speakers, we can easily tell which word comes from which speaker confidently without using any acoustic speaker diarization system. In practice, diarization errors can be much more complicated than the simple example in Fig. 1. To handle such cases, we propose DiarizationLM, a framework to post-process the orchestrated ASR and speaker diarization outputs with a large language model (LLM). While the experiments performed in this paper mainly focus on reducing word diarization errors using LLM, we also show examples how LLM can be used for other purposes such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module in Appendix A.
@@ 3.4 Transcript-Preserving Speaker Transfer
Below we show a simple example of the inputs and output of the TPST algorithm:
Source words: hello good morning hi how are you pretty good
Source speakers: 1 1 1 2 2 2 2 1 1
Target words: hello morning hi hey are you be good
Target speakers: 1 2 2 2 1 1 2 1
Transfered speakers: 1 1 2 2 2 2 1 1
@@ 4.3 Models
For the ASR model in Fig. 2, we use a universal speech model (USM)  [44] with 600 million parameters trained with the RNN-T loss [45]. For the speaker diarization model in Fig. 2, we use the turn-to-diarize system [7] with a multi-stage clustering setup [17] in our experiments, which is capable of diarizing hours of audio recordings in real time on a mobile device [24]. The number of speakers is unknown (from $1$ to $\infty$) to the speaker diarization system in all of our experiments. However, we would like to point out that the proposed framework is very generic and should work with other ASR or speaker diarization systems as well, such as variants of end-to-end speaker diarization models [20, 21, 22, 23].
@@
For the LLM in Fig. 2, we experiment with the PaLM 2-S model (“text-bison” model in Google Cloud API) and the PaLM 2-L model (“text-unicorn” model in Google Cloud API) [46]. We use the PaLM 2-S model as our foundation model, and finetune it on the dataset described in Section 4.1 with data processing steps described in Section 3.5. This model uses a sentence piece model (SPM) of 256k tokens as its tokenizer [47]. During finetuning, we limit the LLM input size by 4,096 tokens, and segment our training and testing data accordingly. The PaLM 2-L model will only be used for zero-shot and one-shot experiments, as described in Section 4.4.
@@
In our prompt builder module, we use an empty prompt prefix, and a 5-character prompt suffix “ --> ” (note the two spaces around the arrow). For the completions in our training data, we use a 6-character completion suffix “ [eod]” (short for “end of document”; note the leading space). After processing the training data with the prompt builder module, we result in 13,430 prompt-completion pairs for training in total. The average length of a prompt is 2,371 SPM tokens, and the average length of a completion is 2,329 tokens. The LLM is trained for 1,200 steps with a batch size of 16.
4.4 Zero-shot and one-shot baselines
Apart from finetuning the PaLM 2-S model on the speaker diarization task, we also experiment with directly using the PaLM 2-S and PaLM 2-L models on the speaker diarization task without finetuning. This is more similar to the example we demonstrated in Section 2.
For the zero-shot setup, we use a prompt prefix that contains an instruction describing the task, as shown below.
Prompt prefix for zero-shot:
In the speaker diarization transcript below, some words are potentially misplaced.
Please correct those words and move them to the right speaker.
[…]
For the one-shot setup, the prompt prefix contains both the instruction describing the task, and also a small example, as shown below.
Prompt prefix for one-shot:
In the speaker diarization transcript below, some words are potentially misplaced. Please correct those words and move them to the right speaker. For example, given this input transcript,
<spk:1> How are you doing today? I <spk:2> am doing very well. How was everything at the <spk:1> party? Oh, the party? It was awesome. We had lots of fun. Good <spk:2> to hear!
The correct output transcript should be:
<spk:1> How are you doing today? <spk:2> I am doing very well. How was everything at the party? <spk:1> Oh, the party? It was awesome. We had lots of fun. <spk:2> Good to hear!
[…]
4.5 Evaluation results
In Table 1, we show the evaluation results of the USM + turn-to-diarize baseline together with the outputs post-processed by DiarizationLM. We report results for zero-shot, one-shot, and finetuning on the diarization task with three different flavors.
For zero-shot and one-shot experiments with PaLM 2-S, we observe significantly worse WDER and cpWER performance compared with the baseline system, indicating the PaLM 2-S foundation model does not offer speaker diarization capabilities without finetuning. Zero-shot experiment with PaLM 2-L model also shows bad performance, while one-shot experiment with PaLM 2-L model is much better, but still worse than the baseline system. Our results indicate that the PaLM 2-L model with one-shot is able to improve speaker diarization in relatively simple cases as shown in Section 2 and Appendix A. However, real world applications can be much more complicated with errors from both the ASR system and the speaker diarization system. In such cases, even with one-shot, LLM can still introduce even more errors to the results if not finetuned specifically on the speaker diarization task.
@@ 4.5 Evaluation results
On both datasets, we observe big improvement of both WDER and cpWER with any of the three finetuning flavors. Interesting, the biggest improvement is observed with the hyp2ora flavor, while the smallest improvement is observed with the deg2ref flavor. Specifically for hyp2ora, we see a rel. 55.5% improvement of WDER after post-processing with DiarizationLM on the Fisher testing set. Even if we did not use any Callhome data during the LLM finetuning, we see a rel. 44.9% improvement of WDER on the Callhome testing set. The WER of the USM on the two testing sets are relatively high due to domain mismatch and suboptimal annotation quality of the ground truth. However, this also demonstrated that the DiarizationLM solution provides consistent quality gains even with out-of-domain ASR and speaker diarization models.
@@ 4.6 Case studies
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. WERs are the same for all systems due to TPST. All numbers are percentages.
System Fisher testing set Callhome testing set
WER WDER cpWER WER WDER cpWER
USM + turn-to-diarize baseline 15.48 5.32 21.19 15.36 7.72 24.39
+ PaLM 2-S zero-shot - 11.96 30.19 - 12.26 30.60
+ PaLM 2-S one-shot - 16.58 38.03 - 14.50 34.32
+ PaLM 2-L zero-shot - 11.36 31.78 - 13.29 34.30
+ PaLM 2-L one-shot - 5.94 22.21 - 7.95 24.67
+ PaLM 2-S finetuned (hyp2ora flavor) - 2.37 16.93 - 4.25 20.22
+ PaLM 2-S finetuned (deg2ref flavor) - 3.94 18.55 - 5.33 21.47
+ PaLM 2-S finetuned (mixed flavor) - 2.41 16.94 - 4.76 20.84
Table 2: Evaluation results of the turn-to-diarize baseline system with reference ASR transcript (assuming WER=0%) and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. All numbers are percentages.
System Fisher testing set Callhome testing set
@@
Reference + turn-to-diarize baseline 2.81 5.19 3.74 6.82
+ PaLM 2-S zero-shot 7.50 12.70 7.29 12.79
+ PaLM 2-S one-shot 10.92 19.16 12.79 21.65
+ PaLM 2-L zero-shot 8.69 16.85 11.67 22.87
+ PaLM 2-L one-shot 3.23 5.99 3.76 6.95
+ PaLM 2-S finetuned 1.18 2.21 1.49 2.66
We also look into why zero-shot and one-shot experiments in Table 1 produced worse results than the baseline system. We found that without finetuning on the speaker diarization tasks, zero-shot and one-shot outputs from the LLM often delete big chunks of hypothesis text from the prompt. Finetuning the LLM is critical to avoid such undesired deletions. A few zero-shot examples with the PaLM 2-S model from the Fisher testing set were shown in Table 5.
@@
Table 5: Example cases from the Fisher testing set where zero-shot PaLM 2-S deletes lots of text from the prompt.
Utterance Before DiarizationLM After DiarizationLM
fe_03_11252
<spk:1> oh okay i believe it’s a lot wrong with the public schools i don’t believe that they’re um that they’re giving um these kids a sense of um well they’re not teaching them what they need to know once they get out of um school you know mhm um what what’s happening is that’s probably why you got a lot of um a lot of people that’s unemployed i think you know they you get a lot from school and they taking a lot of um i guess the economics out of school you know
<spk:2> right
<spk:1> oh okay i believe it’s a lot wrong with the public schools i don’t believe that they’re um that they’re giving um these kids a sense of um well they’re not teaching them what they need to know once they get out of um school you know
<spk:2> right
fe_03_11224
<spk:1> so um i think what do you think is an important thing in a relation i think the topic was um what you um what are the most important things in a life partner yeah uh h well what do you think me
<spk:2> i would have to say trust and honesty like cuz without that you really don’t have nothing to build on you know right yeah
<spk:1> so um i think what do you think is an important thing in a relation
<spk:2> i would have to say trust and honesty like cuz without that you really don’t have nothing to build on you know right
@@ 5 Discussion and future work
Another research direction would be to compare different LLMs, in different size variants on the speaker diarization task. Specifically, the performance will likely be even better if we finetune larger models such as PaLM 2-M or PaLM 2-L. It would also be interesting to reproduce the experiments with other speaker diarization systems such as EEND [20] or WEEND [34].